Solve the system of linear congruences: 7x = 1 mod 8, 7x = 1 mod 13, 7x = 1 mod 17
Answer #1

7x=1 mod 8

-x=1 mod 8

x=-1=1 mod 8

x=8k+1

7(8k+1)=1 mod 13

56k=-6 mod 13

28k=-3 mod 13

2k=-3 mod 13

Multiplying by 6 gives

12k=-6 mod 13

-k=-6 mod 13

k=6 mod 13

k=13p+6

x=8k+1=104p+49

7x=1 mod 17

We first convert coefficinet of x to 1

5*7x=5 mod 17

x=5 mod 17

104p+49=5 mod 17

2p=-44 mod 17

2p=-51+7=7 mod 17

2p=7 mod 17

Multiplying by 9

18p=p=63=12 mod 17

p=12 mod 17

p=17q+12

x=104p+49=104*17q+104*12+49

x=1768q+1297